OpenStudy (anonymous):

a car travels along a straight line for first half time with speed 40km/h and the second half time with speed 60km/h.find the average speed of car

OpenStudy (neer2890):

240/5=48 km/hr

OpenStudy (neer2890):

Average speed= total distance covered / total time taken so, avg speed=\[\frac{ D/2+D/2 }{ D/2*40 +D/2*60}=240/5=48\]